When calcium carbonate is heated ,it decomposes to produce quicklime and carbon dioxide




Calculate the volume of carbon dioxide produced at STP from the decomposition of 76g of CaCO3

1
Expert's answer
2022-05-05T16:32:02-0400

CaCO3=CaO+CO2, n(CaO)=n(CO2)=76/100=0.76 mol, V(CO2)=22.4*0.76=17.024 liters